# JEPA4Rec: Learning Effective Language Representations for Sequential Recommendation via Joint Embedding Predictive Architecture

### v1

@@ Abstract
Language representation learning has emerged as a promising approach for sequential recommendation, thanks to its ability to learn generalizable representations. However, despite its advantages, this approach still struggles with data sparsity and a limited understanding of common-sense user preferences. To address these limitations, we propose JEPA4Rec, a framework that combines Joint Embedding Predictive Architecture with language modeling of item textual descriptions. JEPA4Rec captures semantically rich and transferable representations, improving recommendation performance and reducing reliance on large-scale pre-training data. Specifically, JEPA4Rec represents items as text sentences by flattening descriptive information such as title, category, and other attributes. To encode these sentences, we employ a bidirectional Transformer encoder with modified embedding layers tailored for capturing item information in recommendation datasets. We apply masking to text sentences and use them to predict the representations of the unmasked sentences, helping the model learn generalizable item embeddings. To further improve recommendation performance and language understanding, we employ a two-stage training strategy incorporating self-supervised learning losses. Experiments on six real-world datasets demonstrate that JEPA4Rec consistently outperforms state-of-the-art methods, particularly in cross-domain, cross-platform, and low-resource scenarios.
@@ 1 Introduction
Sequential recommendation predicts the next item a user will interact with based on past behavior. Traditional ID-based methods capture sequential patterns well but struggle with cold-start items and knowledge transfer across domains (Fang et al., 2020; Kang & McAuley, 2018; Sun et al., 2019). To address this issue, cross-domain methods leverage overlapping users or items (Tang et al., 2012; Zhu et al., 2021b), however their real-world applicability is limited due to the scarcity of shared data. Another approach utilizes modalities such as text or images, but the semantic gap between domains remains a challenge (Yuan et al., 2023). Pretrained language models (PLMs), trained on general data like Wikipedia (Devlin et al., 2019), often fail to align with item descriptions and generate embeddings at the sentence level, limiting their ability to effectively model user preferences (Liu et al., 2023).
@@
Building upon this, we propose JEPA4Rec, a framework that leverages JEPA for language representation learning in sequential recommendation. To enrich the semantic meaning of item representations, we transform item metadata (e.g., title, category, and description) into a single text sentence (illustration in Figure 1). Effectively learning item sentence representations and capturing common sense in user preferences is crucial for recommendation. Here, common sense refers to the model’s ability to grasp inherent user behaviors, such as brand loyalty and category preferences, beyond explicit interaction patterns. This is where JEPA plays a central role, enabling structured and transferable representation learning. Our main contributions are as follows:
@@
    We construct a bidirectional Transformer encoder with modified embedding layers tailored for encoding item sentences, which are then used as both the Context and Target Encoders. Additionally, we employ tokens masking strategy that selectively hide history item information at varying rates. This requires the model to reconstruct missing details from partial item information, allowing generalizable item representations and enhancing common-sense preference learning.
@@
    We adopt a two-stage training approach (pre-training and finetuning) leveraging self-supervised learning objectives tailored for both recommendation and language understanding tasks.
@@ 2 Related work
Sequential Recommendation aims to predict users’ next interactions by modeling historical behaviors. Traditional methods, including RNNs (Hidasi et al., 2015; Li et al., 2017), CNNs (Tang & Wang, 2018), and Transformers (Sun et al., 2019; Assran et al., 2023), rely on item IDs, limiting transferability. Recent works incorporate textual features (Hou et al., 2022; 2023) but separate language understanding from sequential modeling. Our work unifies these aspects, leveraging JEPA for semantically rich, transferable item representations.
Transfer Learning for Recommendation addresses data sparsity by leveraging shared knowledge (Singh, 2020). PLMs generate universal item representations (Devlin et al., 2019; Geng et al., 2022) but require large-scale pre-training (Liu et al., 2023) and tightly couple text and item representations (Hou et al., 2023; 2022). JEPA4Rec overcomes these issues by treating item metadata as text, using a Transformer with tailored embeddings and JEPA-based pre-training to enhance generalization, reduce pre-training dependency, and improve cross-domain performance without relying on common users (Zhu et al., 2021a).
@@ 3 Methodology
Figure 2: Overview of JEPA4Rec. Left: The items are represented as text sentences. The model leverages the user’s historical interactions along with partial text information to predict the full representations of the items. Specifically, the historical item sequence embedding $\mathbf{h}_{CLS}$ is combined with the item position embedding $\mathbf{D}_{n}$ of item $i_{n}$ and the embedding of the next item $\mathbf{h}_{n+1,M}$ when they are masked. These representations are then passed through the predictor to reconstruct the complete representations of $i_{n}$ and $i_{n+1}$, which are obtained from the Target Encoder. Right: The pre-training process enhances language understanding using Masked Language Modeling, $L_{2}$ loss for generalizable embeddings, and contrastive learning to optimize recommendation accuracy.
@@ 3.2 Item representation
Encode Item Mechanism To encode item sentences, we follow previous work in Sun et al. (2019); Geng et al. (2022) to construct four different types of embeddings:
@@
where $l$ is the maximum sequence length. To encode $\mathbf{E}_{X}$, we use Longformer (Beltagy et al., 2020), a bidirectional Transformer optimized for long sequences. Its local windowed attention mechanism enables efficient encoding of $X$. Similar to Longformer’s document processing setup, the special token $[CLS]$ has global attention, while other tokens rely on local windowed attention. The model generates $d$-dimensional word representations as follows:
@@
where each word representation $\mathbf{h}_{w}\in\mathbb{R}^{d}$. Following standard language model practices, we use $\mathbf{h}_{CLS}$ as the sequence representation. Although we use Longformer, other Transformer encoders such as BERT (Devlin et al., 2019) or RoBERTa (Liu et al., 2019) can also be applied for encoding. To encode items, JEPA4Rec treats each item as a single-item sequence $X=\{[CLS],S_{i}\}$ and obtains its embedding $\mathbf{h}_{i}$ from the sequence representation. We use the same embedding layer and Longformer encoder for both the Context Encoder $f_{\theta}$ and Target Encoder $f_{\overline{\theta}}$, with the latter updated via an exponential moving average for stable training and prevent representation collapse (Chen & He, 2021).
@@ 3.3 Pre-training Framework
The goal of pre-training is to establish a strong parameter initialization for downstream tasks. Our approach integrates both language understanding and recommendation learning and enhances the model’s ability to learn common sense in user preferences, improve item representation, and create generalizable item embeddings without relying on large-scale pre-training data. Therefore, JEPA4Rec is pre-trained using three key objectives: Masked Language Modeling (MLM), Mapping Representation task and Sequence-Item contrastive task.
@@ 3.3.1 Masking strategy
Since we want the model to infer $i_{n+1}$ based on the user’s history and partial information from $S_{n+1}$, we mask a significant portion ($50\%$) of tokens in $S_{n+1}$ and train the model to predict the full representation of $i_{n+1}$. For history sentences, excessive masking could degrade the History Sequence Embedding $\mathbf{h}_{CLS}$, leading to information loss. Therefore, we apply a $15\%$ token masking rate, consistent with pre-trained language model studies. For both history sequences and the next item, we apply the same masking strategy in BERT: (1) replacing tokens with [MASK] ($80\%$), (2) replacing with a random token ($10\%$), and (3) keeping the original token unchanged ($10\%$). Following this principle, after masking tokens in both history item sentences and the next item, we utilize the history sequence embedding $\mathbf{h}_{CLS}$ to predict their full representations. However, since the number of masked tokens varies across user sequences, the number of masked items within each sequence also differs. To simplify this process, we sample one masked item per sequence and introduce a learnable zero vector when no tokens are masked.
@@ 3.3.2 Learning Framework
JEPA4Rec integrates Mapping Representation task into the training process to enhance item representation learning. To achieve this, we propose a lightweight MLP-based Predictor, which utilize the history sequence embedding $\mathbf{h}_{CLS}$ along with the Item Position Embedding $\mathbf{D}_{n}$, which corresponds to masked history items. This design enables the model to predict the full representation of $S_{i}$, mitigating information loss in $\mathbf{h}_{CLS}$ when encoding long sequences. The predicted representation is computed as:
@@
$\mathcal{L}_{S-I}=-log\dfrac{e^{\operatorname{sim}(\mathbf{h}_{CLS},\mathbf{h}_{i}^{+})/\tau}}{\sum_{i\in B}e^{\operatorname{sim}(\mathbf{h}_{CLS},\mathbf{h}_{i})/\tau}}$
where $\operatorname{sim}$ is the cosine similarity score between 2 vectors; $\mathbf{h}_{i}^{+}$is the representation of the ground truth next item; $\mathcal{B}$ is the ground truth item set in one batch; and $\tau$ is a temperature parameter. At the pre-training stage, we use a multi-task training strategy to jointly optimize JEPA4Rec:
@@
where $\lambda_{1},\lambda_{2}$ is a hyper-parameter to control the weight of MLM and S-I task loss. The pre-trained model will be finetuned for new target domains.
@@ 3.4 Finetuning Framework
After the pre-training phase, the model is capable of zero-shot recommendation on the target domain. We encode all item sentences to construct a dynamic, learnable item matrix $\mathcal{I}$, which enables probability computation for next-item prediction over the entire dataset:
@@
$\mathcal{L}_{FT}=-\log\frac{e^{\operatorname{sim}\left(\mathbf{h}_{CLS},\mathbf{h}_{i}^{+}\right)/\tau}}{\sum_{i\in I}e^{\operatorname{sim}\left(\mathbf{h}_{CLS},\mathbf{h}_{i}\right)/\tau}}$
@@ 3.5 Discussion
In this section, we compare JEPA4Rec with prior sequential recommendation methods to highlight our key innovations. Traditional models such as SASRec Kang & McAuley (2018), and BERT4Rec (Sun et al., 2019) rely on item IDs and trainable embeddings, making them susceptible to data sparsity and ineffective for cold-start items. To address this, context-aware models like UniSRec (Hou et al., 2022), S3-Rec (Zhou et al., 2020), and VQ-Rec (Hou et al., 2023) incorporate side information (e.g., categories, titles) by extracting item features from language models before fusing them into independent sequential models.
Unlike these approaches, JEPA4Rec encodes item sentences using a bidirectional Transformer encoder, but introduces two major improvements: (1) a novel masking strategy that enables learning item representations directly in the embedding space rather than only at the token level, and (2) a joint embedding predictive framework that refines item representations through self-supervised learning. By integrating these innovations, JEPA4Rec enhances common-sense learning in user preferences and improves the generalizability of item embeddings, leading to better adaptation in new domains and more effective cold-start recommendations.
@@ 4 Experiments
In this section, we first set up the experiments and then empirically demonstrate the effectiveness of our proposed model, JEPA4Rec.
@@ 4.1 Experimental Setup
Datasets To evaluate JEPA4Rec’s performance, we conduct pre-training and finetuning using various Amazon review datasets (Hou et al., 2024). The dataset statistics after preprocessing are presented in Table 1. For pre-training, we utilize data from only three categories: Automotive, Grocery and Gourmet Food, and Movies and TV, accounting for approximately $35\%$ of the dataset size used in prior studies (Li et al., 2023a; Hou et al., 2022; 2023). These categories serve as the source domain datasets.
@@
Baselines We compare JEPA4Rec against three categories of baseline models: (1) ID-only Methods: SASRec (Kang & McAuley, 2018), BERT4Rec (Sun et al., 2019); (2) ID-text Methods: S3-Rec (Zhou et al., 2020), LlamaRec; (3) Text-only Methods: UniSRec (Hou et al., 2022), VQ-Rec (Hou et al., 2023), RecFormer (Li et al., 2023a). Detailed descriptions of these models can be found in the Appendix B.
Evaluation Settings We use NDCG@10, Recall@10, and MRR as metrics, applying a leave-one-out strategy: the latest interaction for testing, the second latest for validation, and the rest for training. The ground-truth item is ranked among all items, and average scores are reported. To ensure a fair comparison with RecFormer, the state-of-the-art method, we adopt the same experimental settings as RecFormer, detailed in Appendix C. Other baselines follow prior work settings.

### v3

# Learning Effective Language Representations for Sequential Recommendation via Joint Embedding Predictive Architecture

@@ Abstract
Language representation learning has emerged as a promising approach for sequential recommendation, thanks to its ability to learn generalizable representations. However, despite its advantages, this approach still struggles with data sparsity and a limited understanding of common-sense user preferences. To address these limitations, we propose JEPA4Rec, a framework that combines Joint Embedding Predictive Architecture with language modeling of item textual descriptions. JEPA4Rec captures semantically rich and transferable representations, improving recommendation performance and reducing reliance on large-scale pre-training data. Specifically, JEPA4Rec represents items as text sentences by flattening descriptive information such as title, category, and other attributes. To encode these sentences, we utilize a bidirectional Transformer encoder with modified embedding layers specifically designed to capture item information in recommendation datasets. We apply masking to text sentences and use them to predict the representations of the unmasked sentences, helping the model learn generalizable item embeddings. To further enhance recommendation performance and language understanding, we employ a two-stage training strategy that incorporates self-supervised learning losses. Experiments on six real-world datasets demonstrate that JEPA4Rec consistently outperforms state-of-the-art methods, particularly in cross-domain, cross-platform, and low-resource scenarios.
@@ 1 Introduction
Sequential recommendation predicts the next item a user is likely to interact with based on their past behavior. Traditional ID-based methods capture sequential patterns well but struggle with cold-start items and knowledge transfer across domains (Fang et al., 2020; Kang & McAuley, 2018; Sun et al., 2019). To address this issue, cross-domain methods leverage overlapping users or items (Tang et al., 2012; Zhu et al., 2021b), however, their real-world applicability is limited due to the scarcity of shared data. Another approach utilizes modalities such as text or images, but the semantic gap between domains remains a challenge (Yuan et al., 2023). Pretrained language models (PLMs), trained on general data like Wikipedia (Devlin et al., 2019), often fail to align with item descriptions and generate embeddings at the sentence level, limiting their ability to effectively model user preferences (Liu et al., 2023).
@@
Building upon this, we propose JEPA4Rec, a framework that leverages JEPA for language representation learning in sequential recommendation. To enrich the semantic meaning of item representations, we transform item metadata (e.g., title, category, and description) into a single text sentence (illustration in Figure 1). Effectively learning item sentence representations and capturing common sense in user preferences is crucial for recommendations. Here, common sense refers to the model’s ability to grasp inherent user behaviors, such as brand loyalty and category preferences, beyond explicit interaction patterns. This is where JEPA plays a central role, enabling structured and transferable representation learning. Our main contributions are as follows:
@@
    We construct a bidirectional Transformer encoder with modified embedding layers tailored for encoding item sentences, which are then used as both the Context and Target Encoders. Additionally, we employ a tokens masking strategy that selectively hides history item information at varying rates. This requires the model to reconstruct missing details from partial item information, allowing generalizable item representations and enhancing common-sense preference learning.
@@
    We adopt a two-stage training approach (pre-training and fine-tuning), leveraging self-supervised learning objectives tailored for both recommendation and language understanding tasks.
@@ 2 Related work
Sequential Recommendation aims to predict users’ next interactions by modeling historical behaviors. Traditional methods, including RNNs (Hidasi et al., 2015; Li et al., 2017), CNNs (Tang & Wang, 2018), and Transformers (Sun et al., 2019; Assran et al., 2023), rely on item IDs, limiting transferability. A key limitation of this research direction is its reliance on discrete item IDs, which are inherently non-transferable across domains and incapable of capturing the rich semantic information embedded in item metadata. In contrast, textual item representations offer a more generalizable and semantically meaningful alternative. Approaches like (Hou et al., 2022; Geng et al., 2022) use PLMs to embed item descriptions, allowing knowledge to transfer across platforms and domains. However, these methods often separate representation learning from user modeling, limiting their flexibility. JEPA4Rec addresses this gap by jointly learning from item metadata in natural language form, enabling rich semantic understanding and transfer learning without reliance on shared item IDs or overlapping users (Zhu et al., 2021a; Tang et al., 2012). This makes it particularly suitable for cold-start or low-resource scenarios.
Transfer Learning for Recommendation addresses data sparsity by leveraging shared knowledge (Singh, 2020). PLMs generate universal item representations (Devlin et al., 2019; Geng et al., 2022) but require large-scale pre-training (Liu et al., 2023) and tightly couple text and item representations (Hou et al., 2023; 2022). Unlike contrastive loss, which encourages instance discrimination, or MLM, which focuses on local token-level semantics, JEPA4Rec introduces a novel L2 Mapping Loss that directly aligns user behavior embeddings with item embeddings. This predictive alignment enables the model to reason in the embedding space rather than over discrete tokens, allowing it to capture user preferences more effectively. Through this architecture, JEPA4Rec can reconstruct complete item semantics from partial textual cues, such as fragments of product descriptions or missing metadata. Furthermore, the model can learn representations that generalize across different domains, even when pre-training data is limited. This embedding-level learning paradigm facilitates better user intent modeling, improves robustness to sparse input, and enhances cross-domain performance without requiring overlapping users or shared item identifiers.
@@ 3 Methodology
Figure 2: Overview of JEPA4Rec. Left: The items are represented as text sentences. The model leverages the user’s historical interactions along with partial text information to predict the full representations of the items. Specifically, the historical item sequence embedding $\mathbf{h}_{CLS}$ is combined with the item position embedding $\mathbf{D}_{n}$ of item $i_{n}$ and the embedding of the next item $\mathbf{h}_{n+1,M}$ when they are masked. These representations are then passed through the predictor to reconstruct the complete representations of $i_{n}$ and $i_{n+1}$, which are obtained from the Target Encoder. Right: JEPA4Rec includes the Context Encoder ($f_{\theta}$), which encodes masked item sequences; the Target Encoder ($f_{\overline{\theta}}$), which processes the full item information and is updated via Exponential Moving Average (EMA) of the Context Encoder; and the Predictor ($g_{\phi}$), which aims to recover the complete item representations from the partially observed inputs.
@@ 3.2 Item representation
Encode Item Mechanism To encode item sentences, we follow previous works (Sun et al., 2019; Geng et al., 2022) and construct four types of embeddings. First, token embeddings $\mathbf{A}\in\mathbb{R}^{V_{w}\times d}$ represent the meaning of each word token, where $V_{w}$ is the vocabulary size and $d$ is the embedding dimension. Unlike ID-based approaches (Hua et al., 2023), JEPA4Rec uses text tokens, making the representation more flexible and domain-independent. Second, token position embeddings $\mathbf{B}\in\mathbb{R}^{d}$ capture the position of each token in a sentence, helping the model learn word order. Third, token type embeddings $\mathbf{C}_{CLS},\mathbf{C}_{\text{Attribute}},\mathbf{C}_{\text{Value}}\in\mathbb{R}^{d}$ indicate whether a token belongs to the [CLS] token, an attribute key, or an attribute value, which is crucial for handling repeated structures in item metadata. Finally, item position embeddings $\mathbf{D}\in\mathbb{R}^{n\times d}$ encode the position of items in a user’s interaction sequence, where each item $S_{i}$ shares the same vector $\mathbf{D}_{i}$, aligning all tokens within an item and facilitating sequence modeling. Details about the tokenizer can be found in Appendix B.
@@
Where $l$ is the maximum sequence length. To encode $\mathbf{E}_{X}$, we use Longformer (Beltagy et al., 2020), a bidirectional Transformer optimized for long sequences. Similar to Longformer’s document processing setup, the special token $[CLS]$ has global attention, while other tokens rely on local windowed attention. The model generates $d$-dimensional word representations as follows:
@@
where each word representation $\mathbf{h}_{w}\in\mathbb{R}^{d}$. Following standard language model practices, we use $\mathbf{h}_{CLS}$ as the sequence representation. To encode items, JEPA4Rec treats each item as a single-item sequence $X=\{[CLS],S_{i}\}$ and obtains its embedding $\mathbf{h}_{i}$ from the sequence representation. We use the same embedding layer and Longformer encoder for both the Context Encoder $f_{\theta}$ and Target Encoder $f_{\overline{\theta}}$, with the latter updated via an exponential moving average for stable training and to prevent representation collapse (Chen & He, 2021).
@@ 3.3 Pre-training Framework
The goal of pre-training is to establish a strong parameter initialization for downstream tasks. Our approach integrates both language understanding and recommendation learning and enhances the model’s ability to learn common sense in user preferences, improve item representation, and create generalizable item embeddings without relying on large-scale pre-training data. Therefore, JEPA4Rec is pre-trained using three key objectives: Masked Language Modeling (MLM), Mapping Representation, and Sequence-Item contrastive task.
@@ 3.3.1 Masking strategy
Since we want the model to infer $i_{n+1}$ based on the user’s history and partial information from $S_{n+1}$, we mask a significant portion ($50\%$, detailed ablation study is in Appendix G) of tokens in $S_{n+1}$ and train the model to predict the full representation of $i_{n+1}$. For historical sentences, excessive masking could degrade the History Sequence Embedding $\mathbf{h}_{CLS}$, leading to information loss. Therefore, we apply a $15\%$ token masking rate, consistent with pre-trained language model studies. For both history sequences and the next item, we apply the same masking strategy in BERT: (1) replacing tokens with [MASK] ($80\%$), (2) replacing with a random token ($10\%$), and (3) keeping the original token unchanged ($10\%$). Following this principle, after masking tokens in both history item sentences and the next item, we utilize the history sequence embedding $\mathbf{h}_{CLS}$ to predict their full representations. However, since the number of masked tokens varies across user sequences, the number of masked items within each sequence also differs. To simplify this process, we sample one masked item per sequence and introduce a learnable zero vector when no tokens are masked.
@@ 3.3.2 Learning Framework
JEPA4Rec integrates the Mapping Representation task into the training process to enhance item representation learning. To achieve this, we propose a lightweight MLP-based Predictor, which utilizes the history sequence embedding $\mathbf{h}_{CLS}$ along with the Item Position Embedding $\mathbf{D}_{n}$, which corresponds to masked history items. This design enables the model to predict the full representation of $S_{i}$, mitigating information loss in $\mathbf{h}_{CLS}$ when encoding long sequences. The predicted representation is computed as:
@@
$\mathcal{L}_{S-I}=-log\dfrac{e^{\operatorname{sim}(\mathbf{h}_{CLS},\mathbf{h}_{n+1})/\tau}}{\sum_{i\in B}e^{\operatorname{sim}(\mathbf{h}_{CLS},\mathbf{h}_{i})/\tau}}$
where $\operatorname{sim}$ is the cosine similarity score between 2 vectors; $\mathbf{h}_{n+1}$ is the representation of the ground truth next item; $\mathcal{B}$ is the ground truth item set in one batch; and $\tau$ is a temperature parameter. At the pre-training stage, we use a multi-task training strategy to jointly optimize JEPA4Rec:
@@
Where $\lambda_{1},\lambda_{2}$ a hyperparameters to control the weight of the MLM and S-I task loss. The pre-trained model will be fine-tuned for new target domains.
@@ 3.4 Finetuning Framework
After pretraining, only the Context Encoder is used for fine-tuning in the target domain, with the learned user embedding encapsulating rich and transferable user preference information. We encode all item sentences to construct a dynamic, learnable item matrix $\mathcal{I}$, which enables probability computation for next-item prediction over the entire dataset:
@@
$\mathcal{L}_{FT}=-\log\frac{e^{\operatorname{sim}\left(\mathbf{h}_{CLS},\mathbf{h}_{n+1}\right)/\tau}}{\sum_{i\in I}e^{\operatorname{sim}\left(\mathbf{h}_{CLS},\mathbf{h}_{i}\right)/\tau}}$
@@ 3.5 Discussion
We compare JEPA4Rec with prior sequential recommendation methods to highlight its key innovations and advantages. Traditional models such as SASRec (Kang & McAuley, 2018) and BERT4Rec (Sun et al., 2019) rely on trainable item ID embeddings, which are non-transferable across domains and struggle with cold-start problems due to data sparsity. Context-enhanced models like UniSRec (Hou et al., 2022), S3-Rec (Zhou et al., 2020), and VQ-Rec (Hou et al., 2023) attempt to address this by extracting item features from pretrained language models and injecting them into standalone sequential architectures. However, these approaches typically separate textual understanding from sequential modeling and rely heavily on contrastive objectives or side information fusion.
JEPA4Rec addresses these limitations by leveraging JEPA to learn item-level representations directly in the embedding space, rather than at the token level. This enables the model to produce semantically rich and transferable embeddings that better capture user preferences and generalize across domains. Central to our approach is an information-rich user representation, analogous to the $[CLS]$ token in language models, learned through two auxiliary objectives: reconstructing historical item embeddings to preserve long-range preference signals, and recovering full item information from partially masked input to simulate real-world scenarios with incomplete metadata. As a result, JEPA4Rec is well-suited for settings with sparse or partial item descriptions, offering improved adaptability to new domains and cold-start items (see Appendix E for detailed experiments).
@@ 4.1 Experimental Setup
Datasets To evaluate JEPA4Rec’s performance, we conduct pre-training and fine-tuning using various Amazon review datasets (Hou et al., 2024). The dataset statistics after preprocessing are presented in Table 1. For pre-training, we utilize data from only three categories: Automotive, Grocery and Gourmet Food, and Movies and TV, accounting for approximately $35\%$ of the dataset size used in prior studies (Li et al., 2023a; Hou et al., 2022; 2023). These categories serve as the source domain datasets.
@@
Baselines We compare JEPA4Rec against three categories of baseline models: (1) ID-only Methods: SASRec (Kang & McAuley, 2018), BERT4Rec (Sun et al., 2019); (2) ID-text Methods: S3-Rec (Zhou et al., 2020), LlamaRec; (3) Text-only Methods: UniSRec (Hou et al., 2022), VQ-Rec (Hou et al., 2023), RecFormer (Li et al., 2023a). Detailed descriptions of these models can be found in the Appendix C.
Evaluation Settings We use NDCG@10, Recall@10, and MRR as metrics, applying a leave-one-out strategy: the latest interaction for testing, the second latest for validation, and the rest for training. The ground-truth item is ranked among all items, and average scores are reported. To ensure a fair comparison with RecFormer, the state-of-the-art method, we adopt the same experimental settings as RecFormer, detailed in Appendix D. Other baselines follow prior work settings.